# Chinese vs. World Bank Development Projects: Insights from Earth Observation and Computer Vision on Wealth Gains in Africa, 2002-2013

### v1

@@ 3.3 Daytime Satellite Image Confounder Details
Due to clouding and since no single Landsat series covers all study years (Landsat 7 images taken on and after May 31, 2003, also have diagonal lines across them due to a Scan Line Corrector (SLC) hardware failure (Science, 2021, November 30)), we create composite images from the median pixel of all available satellites each period. This per-pixel median mosaic across sensors and dates mitigates the Landsat-7 SLC-off striping by infilling the missing scan lines with valid pixels from overlapping acquisitions. For each cluster, we build fixed, non-overlapping three-year composites—2002–2004, 2005–2007, 2008–2010, and 2011–2013—taking the per-pixel median of all available, cloud-masked Landsat 5/7/8 scenes in the window; these windows are identical across units and do not depend on treatment timing. We use visible light bands (red, green, blue) plus near-infrared and short-wave infrared spectral bands to capture information about built-up areas. More detail is available in the Data Appendix.
@@ 4 Estimation Method: Image-based Confounding Analysis
Part of our analysis relies on computer vision for causal-adjustment methods to process satellite images, developed elsewhere Jerzak (2023a, b). We use Vision Transformers, a computer vision machine learning algorithm, for our primary results, which rely on ViT, because this model is generally considered state-of-the-art (but also reports results for tabular-only, and later, unit fixed effects, analysis).
@@
After assembling the data, we prepare them for modeling. Besides the usual data curation, this preparation entails assigning all the tabular data in the set $\mathbf{X}_{it}$ and the images in the set $\mathbf{M}_{i(\overline{t}-1)}$. The approach here divides each image $i$ into smaller patches for the model to digest efficiently; we use a $16\times 16$ patch size. Then we flatten the patches, preparing them for injection into the model. The tabular data has a different dimensionality, and we must therefore project it to the embedding dimensionality. To this, we also append a summary representation vector called classification token (typically denoted CLS). The data streams are now ready for fusion by concatenating the patch, CLS, and (when used) tabular representations.
@@
Once the transformer finishes learning the class probabilities, we plug them into an Inverse Probability weighting method with the Hájek adjustment, normalizing the treatment propensity scores to sum to one. As shown in the following formulae, the methods used here rely on estimating the ATE by taking a mean difference in the expected value of the post-aid International Wealth Index estimates between neighborhoods that did and did not receive aid, adjusting for pre-project satellite imagery over treatment and control locations, along with tabular data about each location’s treatment status and confounding variables. Formally, this estimation process contains two stages, as follows:
$\hat{\pi}_{it}\;=\;\widehat{P}\bigl(A_{it}=1\mid\mathbf{M}_{i({\overline{t}-1)}},\,X_{i(t-1)}\bigr)\;=\;f\bigl(\mathbf{M}_{i(\overline{t}-1)},\,X_{i(t-1)}\bigr)$ (1)
$\widehat{ATE}\;=\;\frac{1}{T\,n}\sum_{t=1}^{T}\sum_{i=1}^{n}\Biggl(\frac{A_{it}\,\widehat{IWI}_{i(t+1)}}{\hat{\pi}_{it}}\;-\;\frac{(1-A_{it})\,\widehat{IWI}_{i(t+1)}}{1-\hat{\pi}_{it}}\Biggr)$ (2)
@@
  • $\hat{\pi}_{it}=\widehat{P}\bigl(A_{it}=1\mid\mathbf{M}_{i(\overline{t}-1)},\,X_{i(t-1)}\bigr)$ is the estimated treatment probability of neighborhood $i$ at project commitment time.
@@
  • $X_{i(t-1)}$
    are tabular confounders in neighborhood $i$, one lag prior to project commitment, including ADM2 and year fixed effects. These are summarized in Table 2 and in the Tabular Covariates section below.
@@
    is the total number of treated and control observations for the specific funder and sector.
@@
(Note that Hájek adjustment modifies this baseline estimator slightly by normalizing the weights for improved efficiency.
@@ 5.1 Average Treatment Effect by Funder and Sector
Using our ViT, we estimate ATEs for each funder and report four specifications that correspond to Figure 7: (a) a difference-in-means baseline, (b) tabular covariates with ADM2 fixed effects (X+FE), (c1) imagery only (M), and (c2) imagery plus tabular covariates and ADM2 fixed effects (M+X+FE).
Across nearly all sectors, adding imagery (c1 and c2) shifts ATEs downward relative to tabular-only or unadjusted models, indicating that satellite images capture additional confounding information that standard covariates miss. In general, Chinese projects display larger effects than World Bank projects and sectoral heterogeneity. Notably, across sectors, the imagery-only estimates (c1) are similar to the imagery+X+FE estimates (c2), indicating that pre-treatment satellite features absorb most of the selection-relevant variation that ADM2 fixed effects and standard covariates were proxying.
Relative to prior literature on the World Bank, our image-adjusted estimates are more conservative. Under c2, we see clearly positive effects in Education (110), Transport and Storage (210), General Environmental Protection (410), and Trade and Tourism (330). Several other sectors are small or have confidence intervals that include zero—most notably Communications (220), Banking and Financial Services (240), Energy Generation and Supply (230), Agriculture, Forestry and Fishing (310), and especially Industry, Mining, Construction (320). These patterns are broadly consistent with Bitzer and Gören (2018) and Xu et al. (2020) but are generally smaller after image adjustment; in contrast to Bitzer and Gören (2018), we do find statistically significant positive effects for Education under most model specifications.
@@
We next look across all estimation strategies, looking at the maximum and minimum ATE estimates across sector (examining the most conservative estimates across strategies). For the World Bank, the maximum (most conservative) ATE is for sector Trade and Tourism (330), with a value of 6.99; the minimum ATE is for sector Agriculture, Forestry and Fishing (310), with a value of -0.50. For China, the maximum (most conservative) ATE is for sector Emergency Response (700), with a value of 14.32; the minimum absolute ATE is for sector Communications (220), with a value of -3.69.
These sector extremes align with differences in time-to-impact and what the asset-based outcome captures. For the World Bank, Trade & Tourism (330) plausibly operates through market‑access and service‑sector channels that can translate relatively quickly into higher earnings and durable goods (Scott, 2009; Donaldson and Storeygard, 2016). By contrast, sectors like Agriculture (310) or Energy Generation and Supply (230) often entail long build‑out cycles or require household connections before effects materialize in assets, yielding smaller near‑term IWI gains. On the China side, large Emergency Response (700) estimates are consistent with rapid reconstruction, injecting resources and public works (Cavallo and Noy, 2009). These interpretations are descriptive and do not rule out targeting dynamics or residual confounding.
It is also worth briefly dwelling on the few consistently negative estimated treatment effects—namely, World Bank’s Government and Civil Society (150) interventions, along with the Agriculture (310) sector just discussed. Possibly, these negative or near-zero estimates reflect that such projects often target weaker or more institutionally fragile areas, where aid is harder to translate into short-term asset gains, so our conservative image-adjusted models capture selection into more challenging environments rather than genuine harm. That said, both approaches also yield negative effect estimates in the two-way unit-level fixed effects estimates performed in the robustness section (see below), although their confidence intervals also both contain zero there (estimates being -0.12 [-0.35, 0.12] and -0.02 [-0.22, 0.18], respectively). These results, in particular, call for further investigation.
Figure 7: ATEs by sector, funder, and modeling strategy. “FE” refers to ADM2 fixed effects; specifications: (a) Diff‑in‑means, (b) X+FE, (c1) M, (c2) M+X+FE.
@@ 5.2 Explaining Variation in ATE Estimates
To summarize how our estimates move across specifications, we pool all funder–sector results and regress the ATE level on simple indicators for whether a given specification includes tabular covariates ($X$), ADM2 fixed effects (FE), and satellite imagery ($M$), allowing separate intercepts by funder. Table 4 reports two parsimonious models. We report both because Model 1 provides an interpretable marginal‑effects benchmark for $X$, FE, and $M$, while Model 2 relaxes additivity by estimating distinct combination effects (e.g., M+X+FEM{+}X{+}FE), serving as a robustness check that the imagery‑driven shrinkage in ATEs persists.
Several patterns emerge. First, specifications that incorporate imagery ($M$) produce systematically smaller ATEs than otherwise-comparable models that omit imagery. This is the strongest and most stable association in the table and is consistent with our central claim: images proxy the map-based information donors use when allocating projects, improving confounding control and shrinking observational effects toward zero. Second, adding ADM2 fixed effects (FE) is associated with slightly higher relative ATEs on average, whereas including $X$ alone has little systematic association with the ATE once other ingredients are accounted for. Third, these relationships hold when we allow mean differences by funder; the China mean ATE exceeds the World Bank mean ATE in the pooled regressions, echoing the sector-level results above.
Overall fit is high (adjusted $R^{2}$ is 0.83 and 0.82), indicating that these simple model ingredients explain most of the across-specification variation. Any model variant containing $M$ (e.g., $M$, M+XM{+}X, M+FEM{+}\!FE, M+X+FEM{+}X{+}\!FE) sits below the tabular-only benchmark, reinforcing that the downward shift is broad-based rather than driven by a single specification. In short, imagery-based controls consistently make the estimates more conservative, which is exactly what we would expect if they are capturing selection-relevant features that are missing from standard tabular sets.
@@
Funder: China 12.26 (9.56) 13.84 (11.08)
Funder: WB 2.31 (2.97) 3.55 (3.72)
Contains X -1.01 (-0.96)
Contains FE 1.77 (2.02)
Contains M -3.43 (-3.46)
X -0.38 (-0.16)
FE -0.20 (-0.16)
X+FE 1.24 (0.55)
M -4.48 (-4.23)
M + X -7.56 (-4.88)
M + FE -2.03 (-1.52)
M + X +FE -2.30 (-1.30)
@@
Table 4: Outcome: ATE estimates. Estimator: Weighted OLS (inverse variance weights). $t$-values in parentheses, * indicates $p<0.05$.
@@ 5.3 Insights about Aid Assignment Mechanisms
Figure 8 summarizes out‑of‑sample Area Under the Curves (AUC) for predicting which neighborhoods receive projects. AUC values of $0.50$ indicate predictive performance no better than random guessing; $0.70$ means that, in 70% of randomly drawn treated–control pairs, the model ranks the actually treated neighborhood higher; values closer to $1$ indicate highly predictable placement, based on the data and modeling strategy.
Two broad patterns emerge. First, placement is meaningfully non‑random for both funders: models that use satellite imagery alone (M) already separate treated from control units well above the random‑guessing baseline, and adding tabular covariates with ADM2 fixed effects (M+X+FE) yields small, consistent improvements. This supports our core identification strategy—image features proxy the map‑based information officials use when deciding where to place projects.
Figure 8: Out-of-sample Area Under the Curve (AUC), a metric for evaluating model prediction quality, for each funder, sector, and vision backbone. Values closer to one are better. Here, FE refers to ADM2 fixed effects. Out-of-sample observations come from ADM2 areas not used in model training.
Second, there is clear sectoral heterogeneity in predictability. Sectors whose siting is tightly tied to the measured tabular and image covariates are easier to classify, while social sectors show more moderate separability. This is exactly the type of variation we would expect if imagery and tabular variables together capture on‑the‑ground constraints and opportunities that are highly non-uniform across settings.
@@
WB Funder 0.04 (1.55) 0.06 (1.69)
log(N) 0.10 (2.60) 0.09 (2.40)
Class imbalance 0.32 (1.55) 0.49 (1.99)
@@
Intercept 0.81 (43.39) -0.46 (-0.88) -0.57 (-1.09)
Observations 168 168 168
Adjusted R-squared 0.01 0.05 0.06
@@
Table 5: Outcome: Difference in AUC. Baseline funder: China. Estimator: OLS. $t$-values in parentheses, * indicates p<p<0.05.
Table 5 helps quantify apparent differences between funders. In simple comparisons, World Bank sectors modestly look more variable than Chinese sectors. However, the average difference by funder is not statistically distinguishable from zero, and model fit improves. Overall, variability in AUC values is largely compositional, being driven by data volume and treatment/control balance, not intrinsic differences in assignment opacity between World Bank and Chinese funders.
@@
Figure A.I.2 in the Appendix summarizes the relationship between tabular covariates and project placement in the tabular–only model (X+FEX{+}FE) and the image–augmented model (M+X+FEM{+}X{+}FE). As our modeling strategy involves non-linear operations (e.g., Transformers), we cannot simply look at model parameters, as we can with linear models; rather, we need to compute the sensitivity of the model’s predicted treatment probability to changes in each tabular covariate, averaged across all observations, using automatic differentiation (a computational process that efficiently computes the exact derivatives of complex functions by applying the chain rule to a sequence of elementary arithmetic operations).
In these salience figures the yy‑axis lists the tabular variables (one row per covariate) and the xx‑axis lists sectors. Each cell reports a summary of how much a given covariate moves the model’s predicted probability of treatment for that funder–sector. In Figure A.I.2 (Appendix), the values are signed sensitivities: positive values mean that increasing the covariate is associated with a higher predicted chance of a project, negative values mean the opposite. In Figure 9 (here), each cell shows the change in absolute salience when images are added—formally, $\lvert M{+}X{+}FE\rvert-\lvert X{+}FE\rvert$—so red cells indicate that a variable becomes more influential once images are included, blue cells indicate it matters less, and white indicates little change (or too few units for that funder–sector column). Because inputs are standardized before modeling, salience magnitudes are on a comparable scale across variables.
@@
Rather than dwell on each facet, we highlight several observations.
@@
First, adding images mostly shrinks the absolute salience of many widely used spatial proxies—population density, market access (minutes to city), and country–level governance scores—especially for the World Bank (blue panels in Figure 9 are common). This is consistent with our identification claim: imagery carries geospatial structure that standard tabular sets only approximate, so once we condition on images, those variables matter less.
@@
Second, changes for China are more heterogeneous across sectors. In several funder–sector cells, event- or politics‑style variables (e.g., natural disasters, leader birthplace, or UNSC voting) become relatively more salient once images absorb geography (red cells in Figure 9). Qualitatively, this suggests that after we control for what is visible from space, the model reallocates attention toward non‑visual drivers of Chinese siting in specific sectors.
@@
Third, images mainly re-weight variables rather than flip their signs: the direction of association visible in Figure A.I.2 is largely stable before vs. after imaging, but magnitudes compress. Substantively, that means imagery helps separate “where projects can feasibly go” (geography and the built environment) from “why they go there” (political and programmatic factors).
@@
To gauge cross‑funder similarity in these salience profiles, we perform Canonical Correlation Analysis (CCA) on the sector–by–covariate salience matrices from the image‑augmented model. CCA takes two multivariate datasets measured on the same features—in our case, for each sector we have a vector of saliences over tabular variables for the World Bank and another for China—and finds pairs of linear combinations (one per funder) that are maximally correlated across sectors. We treat sectors as observations and covariate saliences as features (columns are standardized so that scale differences do not drive the results). The leading canonical correlation summarizes how similar the two donors’ multivariate emphasis patterns are across sectors: values near 1 imply that the donors load on similar mixtures of covariates; values near 0 imply largely orthogonal emphasis. This method is appropriate here because it compares full profiles without forcing us to pick one covariate at a time, it is invariant to linear rescaling, and it handles collinearity by working with linear combinations rather than individual variables. The leading canonical correlation here is 0.088, indicating limited alignment between Chinese and World Bank siting logics once geography is accounted for. In short: the two donors emphasize different combinations of factors across sectors.
@@
We do not interpret these salience scores causally; they are descriptive of the assignment mechanism under each adjustment set. Their main value here is to show why image adjustment reduces our ATEs: images soak up spatial structure that tabular proxies were previously carrying, yielding more conservative (and, we argue, more credible) effect estimates.
Robustness: Unit Fixed Effects (Within-Neighborhood) Estimates.

### v2

@@ 3.3 Daytime Satellite Image Confounder Details
Due to clouding and since no single Landsat series covers all study years (Landsat 7 images taken on and after May 31, 2003, also have diagonal lines across them due to a Scan Line Corrector (SLC) hardware failure (Science, 2021, November 30)), we create composite images from the median pixel of all available satellites each period. This per-pixel median mosaic across sensors and dates mitigates the Landsat-7 SLC-off striping by infilling the missing scan lines with valid pixels from overlapping acquisitions. For each cluster, we build fixed, non-overlapping three-year composites—1999–2001, 2002–2004, 2005–2007, 2008–2010, and 2011–2013—taking the per-pixel median of all available, cloud-masked Landsat 5/7/8 scenes in the window; these windows are identical across units and do not depend on treatment timing. We use visible light bands (red, green, blue) plus near-infrared and short-wave infrared spectral bands to capture information about built-up areas. More detail is available in the Data Appendix.
@@ 4 Estimation Method: Image-based Confounding Analysis
Part of our analysis relies on a computer‑vision‑based causal-adjustment method for processing satellite images that we developed elsewhere Jerzak (2023a, b). We use Vision Transformers, a computer vision machine learning algorithm, for our primary results, which rely on ViT, because this model is generally considered state-of-the-art (but also reports results for tabular-only, and later, unit fixed effects, analysis).
Relative to this earlier work, our contribution in this paper is to deploy the EO‑ML adjustment framework in an applied setting—a multi‑funder, multi‑sector evaluation—and to add diagnostic layers. In particular, we (i) estimate a set of nested specifications (diff‑in‑means, X+FE, M, M+X+FE) to quantify how much imagery changes estimated ATEs, and (ii) compute assignment‑mechanism diagnostics (out‑of‑sample AUCs and tabular‑covariate salience measures) that characterize how donors condition on visible and tabular features. We treat these steps as methodological validation and refinement rather than as a new estimator.
@@
After assembling the data, we prepare them for modeling. Besides the usual data curation, this preparation entails assigning all the tabular data in the set $\mathbf{X}_{i(t-1)}$ and the images in the set $\mathbf{M}_{i(\overline{t}-1)}$. The approach here divides each image $i$ into smaller patches for the model to digest efficiently; we use a 16 $\times$ 16 patch size. Then we flatten the patches, preparing them for injection into the model. The tabular data has a different dimensionality, and we must therefore project it to the embedding dimensionality. To this, we also append a summary representation vector called classification token (typically denoted CLS). The data streams are now ready for fusion by concatenating the patch, CLS, and (when used) tabular representations.
@@
Once the transformer finishes learning the class probabilities, we plug them into an Inverse Probability weighting method with the Hájek adjustment, normalizing the observation weights to sum to one. As shown in the following formulae, the methods used here rely on estimating the ATE by taking a mean difference in the expected value of the post-aid International Wealth Index estimates between neighborhoods that did and did not receive aid, adjusting for pre-project satellite imagery over treatment and control locations, along with tabular data about each location’s treatment status and confounding variables. Formally, this estimation process contains two stages, as follows:
$\hat{\pi}_{it}\;=\;\widehat{P}\bigl(A_{it}=1\mid\mathbf{M}_{i(\overline{t}-1)},\,\mathbf{X}_{i(t-1)}\bigr)\;=\;f\bigl(\mathbf{M}_{i(\overline{t}-1)},\,\mathbf{X}_{i(t-1)}\bigr)$ (1)
$\widehat{\textrm{ATE}}\;=\;\frac{1}{T\,n}\sum_{t=1}^{T}\sum_{i=1}^{n}\Biggl(\frac{A_{it}\,\widehat{\textrm{IWI}}_{i(t+1)}}{\hat{\pi}_{it}}\;-\;\frac{(1-A_{it})\,\widehat{\textrm{IWI}}_{i(t+1)}}{1-\hat{\pi}_{it}}\Biggr)$ (2)
@@
  • $\hat{\pi}_{it}=\widehat{P}\bigl(A_{it}=1\mid\mathbf{M}_{i(\overline{t}-1)},\,\mathbf{X}_{i(t-1)}\bigr)$ is the estimated treatment probability of neighborhood $i$ at project commitment time.
@@
  • $\mathbf{X}_{i(t-1)}$
    are tabular confounders in neighborhood $i$, one lag prior to project commitment, including ADM2 and year fixed effects. These are summarized in Table LABEL:tab:Descriptives and in the Tabular Covariates section below.
@@
    is the total number of unique neighborhoods for the specific funder and sector.
@@
(Note that Hájek adjustment modifies this baseline estimator slightly by normalizing the weights for improved efficiency (Freedman and Berk, 2008).
@@ 5.1 Average Treatment Effect by Funder and Sector
Using our ViT, we estimate ATEs for each funder and report four specifications that correspond to Figure 7: (a) a difference-in-means baseline, (b) tabular covariates with ADM2 fixed effects (X+FE), (c) imagery only (M), and (d) imagery plus tabular covariates and ADM2 fixed effects (M+X+FE). This subsection also provides our main methodological benchmark: by comparing diff‑in‑means, tabular‑only (X+FE), imagery‑only (M), and imagery‑plus‑tabular (M+X+FE) estimators, we assess how much image‑based confounding control changes estimated ATEs relative to standard approaches.
Because Figure 7 juxtaposes several estimation regimes, a natural question is which one should be treated as the primary causal estimate. Our preferred adjustment set is the fully image-augmented specification (d), which conditions on pre-treatment satellite imagery—intended to proxy the map-based information donors use when selecting project sites—while also incorporating the standard tabular covariates and ADM2 fixed effects that absorb time-invariant local context.
Across many sectors, adding imagery (c and d) shifts ATEs downward relative to tabular-only or unadjusted models, indicating that satellite images capture additional confounding information that standard covariates may miss. In general, Chinese projects display larger effects than World Bank projects and sectoral heterogeneity. Notably, across sectors, the imagery-only estimates (c) are similar to the imagery+X+FE estimates (d), indicating that pre-treatment satellite features absorb most of the selection-relevant variation that ADM2 fixed effects and standard covariates were proxying.
Relative to prior literature on the World Bank, our image-adjusted estimates are more conservative. Under d, we see clearly positive effects in Education (110), Transport and Storage (210), General Environmental Protection (410), and Trade and Tourism (330). Several other sectors are small or have confidence intervals that include or nearly zero—most notably Communications (220), Banking and Financial Services (240), Energy Generation and Supply (230), Agriculture, Forestry and Fishing (310), and especially Industry, Mining, Construction (320). These patterns are broadly consistent with Bitzer and Gören (2018) and Xu et al. (2020) but are generally smaller after image adjustment; in contrast to Bitzer and Gören (2018), we do find positive effects for Education under most model specifications.
@@
We next look across all estimation strategies, looking at the maximum and minimum ATE estimates across sector (taking median across strategies). For the World Bank, the maximum (median) ATE is for sector Trade and Tourism (330), with a value of 12.29; the minimum ATE is for sector Government and Civil Society (150), with a value of -0.16. For China, the maximum median ATE is for sector Emergency Response (700), with a value of 15.15; the minimum is for sector Agriculture, Forestry and Fishing (310), with a value of 1.21.
These sector extremes align with differences in time-to-impact and what the asset-based outcome captures. For the World Bank, Trade & Tourism (330) plausibly operates through market‑access and service‑sector channels that can translate relatively quickly into higher earnings and durable goods (Scott, 2009; Donaldson and Storeygard, 2016). By contrast, sectors like Government and Civil Society (150) or Agriculture, Forestry and Fishing (310) might entail long build‑out cycles or require significant preparatory investments before effects materialize, yielding smaller near‑term IWI gains. On the China side, large Emergency Response (700) estimates are consistent with rapid reconstruction, injecting resources and public works (Cavallo and Noy, 2009). These interpretations are descriptive and do not rule out targeting dynamics or residual confounding.
It is also worth briefly discussing the few consistently negative estimated treatment effects—for example, World Bank’s Government and Civil Society (150) interventions sector just discussed. Possibly, these negative or near-zero estimates might reflect how such projects may target weaker or more institutionally fragile areas, where aid is harder to translate into short-term asset gains, so our conservative image-adjusted models capture selection into more challenging environments rather than genuine harm. That said, the within-neighborhood fixed-effects robustness check in Section 5.4 yields similarly small (and sometimes negative) point estimates with wide confidence intervals for these two sectors, suggesting more limited within-unit evidence of short-run asset gains. These cases merit further investigation.
Figure 7: ATEs by sector, funder, and modeling strategy. “FE" refers to ADM2 fixed effects; specifications: (a) Diff‑in‑means, (b) X+FE, (c) M, (d) M+X+FE.
@@ 5.2 Explaining Variation in ATE Estimates
To summarize how different modeling ingredients affect the estimates, we pool all funder–sector results and regress the ATE level on simple indicators for whether a given specification includes tabular covariates ($X$), ADM2 fixed effects (FE), and satellite imagery ($M$), allowing separate intercepts by funder. This provides a compact diagnostic of the extent to which the EO–ML component contributes beyond conventional tabular adjustment. Table 4 reports two parsimonious models. We report both because Model 1 provides an interpretable marginal‑effects benchmark for $X$, FE, and $M$, while Model 2 relaxes additivity by estimating distinct combination effects (e.g., M+X+FEM{+}X{+}FE), serving as a robustness check that the imagery‑driven shrinkage in ATEs persists.
Several patterns emerge. First, specifications that incorporate imagery ($M$) produce often smaller ATEs than otherwise-comparable models that omit imagery. This is the strongest and most stable association in the table and is consistent with our central claim: images proxy the map-based information donors use when allocating projects, improving confounding control and shrinking observational effects toward zero. Second, adding ADM2 fixed effects (FE) is associated with slightly lower absolute ATEs on average (see robustness section for an individual-level FE analysis), whereas including $X$ alone has little systematic association with the ATE once other ingredients are accounted for. Third, these relationships hold when we allow mean differences by funder; the China mean ATE exceeds the World Bank mean ATE in the pooled regressions, echoing the sector-level results above.
Overall fit is high (adjusted $R^{2}$ is 0.71 and 0.73), indicating that these simple model ingredients explain most of the across-specification variation. Any model variant containing $M$ (e.g., $M$, M+XM{+}X, M+FEM{+}\!FE, M+X+FEM{+}X{+}\!FE) sits below the tabular-only benchmark, reinforcing that the downward shift is broad-based rather than driven by a single specification. In short, imagery-based controls consistently make the estimates more conservative, which is exactly what we would expect if they are capturing selection-relevant features that are missing from standard tabular sets.
@@
Funder: China 13.39 (13.22) 15.66 (11.55)
Funder: WB 6.74 (8.46) 9.05 (7.11)
Contains X -1.55 (-2.15)
Contains FE -1.15 (-1.59)
Contains M -2.75 (-3.84)
X -5.08 (-3.03)
FE -4.92 (-3.25)
X+FE -4.72 (-2.82)
M -6.89 (-4.60)
M + X -6.05 (-3.80)
M + FE -5.35 (-3.42)
M + X +FE -7.30 (-4.85)
@@
Table 4: Outcome: Absolute ATE estimates. Estimator: Weighted OLS (inverse variance weights). $t$-values in parentheses, * indicates $p<0.05$.
@@ 5.3 Insights about Aid Assignment Mechanisms
Having assessed the robustness of our impact estimates, we now turn to the assignment mechanism, using out-of-sample prediction performance to quantify how systematically each donor targets neighborhoods. Figure 8 summarizes out‑of‑sample Area Under the Curves (AUC) for predicting which neighborhoods receive projects. AUC values of 0.50 indicate predictive performance no better than random guessing; 0.70 means that, in 70% of randomly drawn treated–control pairs, the model ranks the actually treated neighborhood higher; values closer to 1 indicate highly predictable placement, based on the data and modeling strategy.
Two broad patterns emerge. First, placement is meaningfully non‑random for both funders: models that use satellite imagery alone (M) already separate treated from control units well above the random‑guessing baseline in most cases, and adding tabular covariates with ADM2 fixed effects (M+X+FE) yields small, consistent improvements. This supports our core identification strategy—image features proxy the map‑based information officials use when deciding where to place projects.
Figure 8: Out-of-sample Area Under the Curve (AUC), a metric for evaluating model prediction quality, for each funder, sector, and analysis backbone. Values closer to one are better. Here, FE refers to ADM2 fixed effects. Out-of-sample observations come from ADM2 areas not used in model training.
Second, there is sectoral heterogeneity in treatment predictability. Projects with physical requirements show often high levels of predictability using image covariates (e.g., mining), while some social sectors show more moderate separability (Education (110) is an exception). This is the type of variation we would expect if imagery and tabular variables together capture on‑the‑ground constraints and opportunities that are highly non-uniform across settings.
@@
WB Funder 0.08 (2.63) 0.08 (2.40)
log(N) 0.17 (4.85) 0.17 (4.81)
Class imbalance 0.40 (2.36) 0.61 (2.97)
@@
Intercept 0.77 (30.42) -1.23 (-2.75) -1.38 (-3.04)
Observations 175 175 175
Adjusted R-squared 0.03 0.18 0.20
@@
Table 5: Outcome: AUC. Baseline funder: China. Estimator: OLS. $t$-values in parentheses, * indicates p<p<0.05.
Table 5 helps quantify apparent differences between funders. In simple comparisons, World Bank sectors modestly look more predictable than Chinese sectors. Overall, variability in AUC values is largely compositional, being driven by data volume and treatment/control balance, and possibly somewhat intrinsic differences in assignment opacity between World Bank and Chinese funders.
@@
Figure A.I.4 in the Appendix summarizes the relationship between tabular covariates and project placement in the tabular–only model (X+FEX{+}FE) and the image–augmented model (M+X+FEM{+}X{+}FE). As our modeling strategy involves non-linear operations (e.g., Transformers), we cannot simply look at model parameters, as we can with linear models; rather, we need to compute the sensitivity of the model’s predicted treatment probability to changes in each tabular covariate, averaged across all observations, using automatic differentiation (a computational process that efficiently computes the exact derivatives of complex functions by applying the chain rule to a sequence of elementary arithmetic operations).
In these salience figures the yy‑axis lists the tabular variables (one row per covariate) and the xx‑axis lists sectors. Each cell reports a summary of how much a given covariate moves the model’s predicted probability of treatment for that funder–sector. In Figure A.I.4 (Appendix), the values are signed sensitivities: positive values mean that increasing the covariate is associated with a higher predicted chance of a project, negative values mean the opposite. In Figure 9 (here), each cell shows the change in absolute salience when images are added, $\lvert\textrm{Salience}(M{+}X{+}FE)\rvert-\lvert\textrm{Salience}(X{+}FE)\rvert$, so red cells indicate that a variable becomes more influential once images are included, blue cells indicate it matters less, and white indicates little change (or too few units for that funder–sector column). Because inputs are standardized before modeling, salience magnitudes are on a comparable scale across variables.
@@
We highlight several observations.
@@
First, adding images often shrinks the absolute salience of many widely used spatial proxies—population density, market access (minutes to city), and country–level governance scores—especially for the World Bank (blue panels in Figure 9 are common). This is consistent with the identification claim: imagery carries geospatial structure that standard tabular sets only approximate, so once we condition on images, those variables matter less.
@@
Second, changes for China are more heterogeneous across sectors. In several funder–sector cells, event- or politics‑style variables (e.g., natural disasters, leader birthplace, or UNSC voting) become relatively more salient once images absorb geography (red cells in Figure 9). Qualitatively, this suggests that after we control for what is visible from space, the model reallocates attention toward non‑visual drivers of Chinese project placement in specific sectors.
@@
Third, images mainly re-weight variables rather than flip their signs: the direction of association visible in Figure A.I.4 is largely stable before vs. after imaging, but magnitudes compress. Substantively, that means imagery helps separate “where projects can feasibly go" (geography and the built environment) from “why they go there" (political and programmatic factors).
@@
To gauge cross‑funder similarity in these salience profiles, we perform Canonical Correlation Analysis (CCA) on the sector–by–covariate salience matrices from the image‑augmented model. CCA takes two multivariate datasets measured on the same features—in our case, for each sector we have a vector of saliences over tabular variables for the World Bank and another for China—and finds pairs of linear combinations (one per funder) that are maximally correlated across sectors. We treat sectors as observations and covariate saliences as features (columns are standardized so that scale differences do not drive the results). The leading canonical correlation summarizes how similar the two donors’ multivariate emphasis patterns are across sectors: values near 1 imply that the donors load on similar mixtures of covariates; values near 0 imply largely orthogonal emphasis. This method is appropriate here because it compares full profiles without forcing us to pick one covariate at a time, it is invariant to linear rescaling, and it handles collinearity by working with linear combinations rather than individual variables. The leading canonical correlation here is 0.182, indicating limited alignment between Chinese and World Bank placement logics once geography is accounted for. In short: the two donors emphasize different combinations of factors across sectors.
@@
We do not interpret these salience scores causally; they are descriptive of the assignment mechanism under each adjustment set. Their main value here is to show why image adjustment reduces ATEs estimates: images soak up spatial structure that tabular proxies were previously carrying, yielding more conservative (and, we argue, more credible) effect estimates.
5.4 Robustness and Alternative Identification Strategies
The main estimates in Figure 7 rely on selection-on-observables after conditioning on pre-treatment imagery and tabular covariates. We run four complementary robustness checks that probe whether the qualitative conclusions depend on (i) identifying variation only within neighborhoods over time, (ii) including lagged nightlights in the covariate set, (iii) the spatial scale of the IWI outcome grid (motivated by DHS coordinate displacement), and (iv) our treatment-footprint rules for imprecisely geocoded projects. Detailed figures and tables are reported in the Results Appendix.
Within-neighborhood (unit) fixed effects.